OpenStudy (anonymous):

What are the methods in solving a quadratic equation by completing the squares?

OpenStudy (anonymous):

ask satellite..it's his fave topic

OpenStudy (anonymous):

very funny

OpenStudy (anonymous):

would you like to do an example? \[x^2+6x-2=0\] \[x^2+6x=2\] \[(x+3)^2=2+3^2=2+9=11\] \[x+3=\pm\sqrt{11}\] \[x=-3\pm\sqrt{11}\]

OpenStudy (anonymous):

http://www.purplemath.com/modules/solvquad3.htm

OpenStudy (anonymous):

if you repeat the method with \[ax^2+bx+c=0\] you get \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Totally forgot that this is how you get the last formula :-) Good to come here